farmer ted goes to the auction each week and buys lots of hay bales. last week he bought 3 truck and 10 wagons full of hay... for a total of 946 bales. this week he bought 9 trucks and 6 wagons full of hay... for a total of 798 bales. how many bales of hay are on each truck, and same for each wagon.
Answer by Boreal(15235)

3T+10H=946
9T+6H=798
-9T-30H=-2838
-24H=-2040
H=85 bales per wagon
T=32 bales per truck